Seawater has a density of 1,025 kg/m3 at the surface. The bulk modulus of seawater is 0.2100 × 1010 N/m2. What is the approximate density of seawater at a depth of 10.00 km?

a)1074 kg/m3

b)1230 kg/m3

c)1574 kg/m3

d)1764 kg/m3

e)1973 kg/m3
Answer #1

volume ration

v'/v = p'/B

B = bulk modulus

v'/v = dgh/B

v'/v = 1025*9.8*10000/2.1*10^9

v'/v = 0.0479

1-v'/v = new volume = 0.9521

d = m/v

density at a depth is = 1025/0.9521= 1076 kg /m^3

or

rho = rho0*exp(p/B)

rho = rho 0exp(1025*9.81*10000/2.1*10^9) = 1075.2 kg/m^3